# iIntel Labs, USA

### v1

@@ Abstract
We present OuroMamba, the first data-free post-training quantization (DFQ) method for vision Mamba-based models (VMMs). We identify two key challenges in enabling DFQ for VMMs, (1) VMM’s recurrent state transitions restricts capturing of long-range interactions and leads to semantically weak synthetic data, (2) VMM activations exhibit dynamic outlier variations across time-steps, rendering existing static PTQ techniques ineffective. To address these challenges, OuroMamba presents a two-stage framework: (1) OuroMamba-Gen to generate semantically rich and meaningful synthetic data. It applies contrastive learning on patch level VMM features generated through neighborhood interactions in the latent state space, (2) OuroMamba-Quant to employ mixed-precision quantization with lightweight dynamic outlier detection during inference. In specific, we present a thresholding based outlier channel selection strategy for activations that gets updated every time-step. Extensive experiments across vision and generative tasks show that our data-free OuroMamba surpasses existing data-driven PTQ techniques, achieving state-of-the-art performance across diverse quantization settings. Additionally, we implement efficient GPU kernels to achieve practical latency speedup of up to $\mathbf{2.36}\times$. Code will be released soon.
Figure 1: Qualitative comparison of VMM implicit attention. (a) Sample input image. (b) Normalized distribution of the input gate values ($\Delta$). Visualization of implicit attention Ali et al. (2024) using (c) the original state ($h$) and (d) the proposed patched state ($h_{p}$), which incorporates spatial dependencies through patched neighborhood interactions.
@@ 1 Introduction
The sub-quadratic compute alternative of the vision State space models (SSMs) Gu & Dao (2023), particularly vision mamba models (VMMs) Zhu et al. (2024) has made them a promising alternative to ViTs Dosovitskiy et al. (2020). Like ViTs, larger VMM variants face deployment challenges due to high memory and latency constraints Li et al. (2025b). Quantization Ramachandran et al. (2024a; b; c) is a popular method to tackle the high memory and latency demand by mapping full-precision (FP) weights and/or activations to lower-bit representations. In particular, post-training quantization (PTQ) Ramachandran et al. (2024a); Kundu et al. (2022) converts a pre-trained FP model to low-precision and requires calibration on a small dataset Ramachandran et al. (2024a) to recover model performance. However, PTQ’s calibration data is typically drawn from the original training dataset, potentially restricting its applicability in situations demanding privacy, security of training set Kundu et al. (2021); Zhang et al. (2023).
@@
Consequently, data-free quantization (DFQ), a subset of post-training quantization (PTQ) has emerged as a promising alternative. DFQ allows calibration via generating synthetic data from Gaussian noise Ramachandran et al. (2024a); Kim et al. (2025); Li et al. (2023b; 2022) mimicking the distribution of the original train-set. In particular, DFQ for ViTs primarily rely on the distinctive response of self-attention to noise and real data to generate synthetic data Li et al. (2023b); Ramachandran et al. (2024a). However, despite their significant progress for ViTs, DFQ techniques for VMM remains largely unexplored. Interestingly, VMM’s recurrent state transition in the S6 layer Gu & Dao (2023) lacks an explicit self-attention mechanism, limiting the usage of any attention based patch similarity improvement. While the S6 layer can be formulated as a variant of linear attention Han et al. (2025), our observations reveal that this implicit attention struggles to distinguish foreground from the background of an image. As observed in Figure 1(c), Figure 4(a), despite VMM’s “theoretical ability” for global interactions Azizi et al. (2025), it fails to capture long-range dependencies in feature representation.
@@
Notably, VMM’s recurrent state transitions introduce dynamic variations in activation characteristics Li et al. (2025b). These variations occur across time steps and tokens [^1] We use token and patch interchangeably based on context., potentially demanding online adaptive outlier selection. This is in sharp contrast to ViTs, which largely demonstrate a static outlier activation pattern Dong et al. (2023). A contemporary work, QMamba Li et al. (2025b), rightly identified the dynamic variations in SSM activations. However, it employs statically determined temporal grouped quantization, unable to adapt to the potential changes in outlier locations, yielding large quantization error at low bit precision. Moreover, QMamba lacks a data-free calibration strategy and limited practical speedup, limiting its use in privacy and latency-sensitive scenarios.
@@
Our Contributions. To address the above challenges in data generation and quantization stages, we present OuroMamba [^2] Inspired by Ouroboros—the self-consuming snake— the name OuroMamba reflects its ability to self-generate data for VMM quantization., a first-of-its-kind method to enable DFQ for VMMs. For data generation, we investigate on VMM’s implicit attention struggles to differentiate foreground with background. In specific, we hypothesize it to be dependent on scanning direction of the compressed hidden state ($h(t)$), which limits explicit spatial token interactions. To mitigate this, we propose OuroMamba-Gen, that essentially enhances the implicit attention through patched spatial interactions in the latent state space, resulting in a refined hidden state namely, patched state $h_{p}(t)$. We then leverage an enhanced implicit attention representation based on $h_{p}(t)$ (Figure 1(d)) to generate synthetic data by employing patch-level contrastive learning Ramachandran et al. (2024a).
@@
For quantization, we first empirically validate that VMM operation leads to dynamic inter-time-step variations of the outlier channel positions. We then present OuroMamba-Quant, a mixed-precision quantization scheme that performs channel-wise static weight quantization and dynamic activation quantization per-time-step during inference. OuroMamba-Quant reduces the quantization error associated to dynamic outliers via three key strategies, (1) a lightweight channel-wise activation outlier detection, (2) mixed-precision quantization, where outlier channels are quantized at higher precision, while remaining inlier channels use lower bit quantization and (3) efficient outlier management using an adaptive outlier list.
To validate the efficacy of OuroMamba, we conduct extensive experimental evaluations on VMMs Zhu et al. (2024); Liu et al. (2025); Huang et al. (2024) and hybrid Transformer-Mamba models Hatamizadeh & Kautz (2025) for classification, detection, segmentation, and generative modeling tasks Hu et al. (2024). OuroMamba achieves SoTA performance over existing data-driven VMM PTQ alternatives, with an accuracy improvement of up to 39%. Additionally, through efficient kernel implementation we demonstrate practical end-to-end latency speedup of up to 2.36$\times$ over FP16 baseline.
@@ 2 Related Works
Data-Driven PTQ. Existing PTQ techniques for ViTs Lin et al. (2021); Jiang et al. (2025); Li et al. (2023c); Yang et al. (2024); Ma et al. (2024); Yuan et al. (2022) address long-tailed distributions and static activation outliers but fail to handle VMMs’ dynamically changing outliers Cho et al. (2025); Li et al. (2025b). While PTQ4VM Cho et al. (2025) adapts SmoothQuant Xiao et al. (2023) to shift activation outliers into weights, it does not quantize SSM activations, limiting its effectiveness. QMamba Li et al. (2025b) employs temporal grouped quantization but relies on static groupings, making it unsuitable for ultra-low bit-widths. OuroMamba is the first to identify dynamic outlier variations in VMMs and introduce a mixed-precision quantization framework that achieves SoTA accuracy at low precision.
Date-Free PTQ. Existing DFQ techniques for ViTs, PSAQ-ViT v1 Li et al. (2022) and v2 Li et al. (2023b) optimize noise into synthetic data by maximizing global patch similarity entropy, but overlook spatial sensitivity and semantic inter-patch relationships, leading to simplistic data. CLAMP-ViT Ramachandran et al. (2024a) addresses this by introducing patch-level contrastive learning on self-attention, aligning foreground patches while separating background patches. Following data generation, these methods apply uniform symmetric quantization to weights and activations, which is suboptimal for VMM. DFQ for VMM thus far remains unexplored, and this work makes the first effort in this direction.
@@ 3.1 Mamba Block in VMM
The S6 Layer. Each Mamba block in VMM utilizes the selective SSM (S6) mechanism Gu & Dao (2023) that maps an input signal $u(t)\in\mathbb{R}^{E}$ to an intermediate hidden state $h(t)\in\mathbb{R}^{E\times N}$ through a linear recurrent transition. Here, $E$,$N,t$ correspond to the number of channels, state dimension and time-step, respectively. $h(t)$ then produces the output $o(t)\in\mathbb{R}^{E}$ via a linear transformation, as shown below.
@@
where, $S^{+}$ is softplus, $\Delta(t)_{\text{proj}},W_{B},\text{ and }W_{C}$ are linear projection layers. $\Delta(t)\in\mathbb{R}^{E\times 1}$ is the discretization tensor. $\bar{A},\bar{B},C,\Delta$ are input-dependent discrete S6 parameters that exhibit dynamic activation variations at each time step. Equation 1 deals with scalar inputs. To operate over an input sequence $u$ of batch size $B$, sequence length $M$ with $E$ channels, $u\in\mathbb{R}^{B\times M\times E}$, Equation 2, Equation 1 are applied independently to each $u(t)\in\mathbb{R}^{B\times E}$ Gu & Dao (2023) across $M$ time-steps.
@@
Selective Scan for Vision. While selective scan in S6 is well-suited for 1D NLP tasks, vision data is 2D, requiring VMMs to adopt specialized scanning strategies to flatten images into 1D sequences before processing them through Mamba’s selective scan. Examples include 2-way scanning Zhu et al. (2024) (Figure 2(a)), 4-way scanning Liu et al. (2025). However, scan direction choice can introduce directional dependency, impacting the model’s ability to capture spatial relationships Xiao et al. (2024).
@@ 3.2 Implicit Attention in VMM
Prior works Han et al. (2025); Ali et al. (2024) have shown that S6 layer operation can be reformulated into an implicit self-attention mechanism. By setting $h(0)=0$ and unrolling Equation 1, the hidden state can be derived as, $h(t)=\sum_{j=1}^{t}(\prod_{k=j+1}^{t}\bar{A}(k))\bar{B}(j)u(j)$ and $o(t)$ can be computed by using this derived value of $h(t)$ in Equation 1. Additionally, in computing $o(t)$, the contribution of the $j^{th}$ token on the $i^{th}$ token is captured as follows,
$\tilde{\alpha}[i,j]=\sum_{m=1}^{N}\alpha^{m}[i,j]=\sum_{m=1}^{N}C(i)(\prod_{k=j+1}^{i}\bar{A}(k)_{m})\bar{B}(j)[m]$ (4)
@@ 3.3 Quantization
In this paper, we perform symmetric group quantization Ramachandran et al. (2024a); Li et al. (2022) of both weights and activations for VMMs. The quantization process for an input FP tensor $X$ with a target bit-precision $b$ is given by,
@@ 3.4 Contrastive Objective
Contrastive learning based on the infoNCE loss Wang et al. (2023) helps learn an anchor patch from both similar (positive) and dissimilar (negative) patches. Following Ramachandran et al. (2024a), the formulation of patch-level contrastive loss ($\mathcal{L}^{C}_{l,(i,j)}$) we employ for data generation is as follows,
@@
where, *, $+$, and -- correspond to the anchor patch, positive, and negative patches, respectively. $\tau$ controls the concentration level Wang et al. (2023). $\lambda_{l,(i,j)}$ (shown without subscript above for brevity) represents the embedding employed to calculate $\mathcal{L}^{C}_{l,(i,j)}$. $l$ is the layer and $(i,j)$ is the location of anchor patch. The final loss is given as $\mathcal{L}^{C}=\sum_{l}\sum_{i}\sum_{j}\mathcal{L}^{C}_{l,(i,j)}$.
@@ 4 Motivational Analysis
Figure 2: W4A8 quantization performance comparison with different calibration data sources.
@@ 4.1 Synthetic Data Generation
We investigate the W4A8 quantization performance of Vim-S Zhu et al. (2024) on an existing VMM PTQ framework, QMamba Li et al. (2025b), using different calibration data sources. In specific, we perform calibration with both real data and ViT generated synthetic data of same sample size of 1024 images. We use synthetic calibration data produced by ViT DFQ technique, namely, CLAMP-ViT Ramachandran et al. (2024a). Figure 2 reveals that across tasks, ViT synthetic data results in significant accuracy degradation (7-14%) as compared to using real calibration data.
@@
As shown in Figure 1(c), VMM’s implicit attention struggles to distinguish foreground from background. This is potentially due to the forgetfulness of the states Azizi et al. (2025) as we traverse along the scanning direction (Figure 4(a)) Han et al. (2025). This, unlike self-attention, limits long-range global interactions. Additionally, VMMs Zhu et al. (2024); Liu et al. (2025) employ multiple scanning directions (subsection 3.1) to improve global interactions. However, they process each direction independently with separate S6 blocks. Consequently, each scanning direction’s implicit attention exhibits directional bias and are incoherent Han et al. (2025).
Figure 4: (a) Forward and backward SSM state transition of Vim-S, (b) Naive synthetic data samples generated by applying Ramachandran et al. (2024a) on VMM implicit attention.
Following Ramachandran et al. (2024a), that applies patch-level contrastive loss to ViT self-attention to generate synthetic data, we apply the same method to VMM’s forward scan S6 block and present the generated data in Figure 4(b). The “naive synthetic data” appears noisy, lacks semantic structure, and performs worse than ViT-generated synthetic data for calibration, Figure 2). Extending this to other scanning directions yields similar results (see section 6). These observations motivate an optimal VMM data generation technique, capturing spatial dependencies without scanning direction constraints.

### v2

@@ Abstract
We present OuroMamba, the first data-free post-training quantization (DFQ) method for vision Mamba-based models (VMMs). We identify two key challenges in enabling DFQ for VMMs, (1) VMM’s recurrent state transitions restricts the capturing of long-range interactions and leads to semantically weak synthetic data, (2) VMM activations exhibit dynamic outlier variations across time-steps, rendering existing static PTQ techniques ineffective. To address these challenges, OuroMamba presents a two-stage framework: (1) OuroMamba-Gen to generate semantically rich and meaningful synthetic data. It applies contrastive learning on patch level VMM features generated through neighborhood interactions in the latent state space, (2) OuroMamba-Quant to employ mixed-precision quantization with lightweight dynamic outlier detection. In specific, we present a thresholding based outlier channel selection strategy for activations that is updated every time-step. Extensive experiments across vision and generative tasks show that our data-free OuroMamba surpasses existing data-driven PTQ techniques, achieving state-of-the-art performance across diverse quantization settings. Additionally, we implement efficient GPU kernels to achieve practical latency speedup of up to $\mathbf{2.36}\times$. Code and synthetic dataset are available here: https://github.com/georgia-tech-synergy-lab/ICCV-OuroMamba.
@@ 1 Introduction
The sub-quadratic compute alternative of the vision State space models (SSMs) [gu2023mamba], particularly vision mamba models (VMMs) [zhu2024vision] has made them a promising alternative to ViTs [dosovitskiy2020image]. Like ViTs, larger VMM variants face deployment challenges due to high memory and latency constraints [li2025qmamba].
@@
Quantization [ramachandran2024clamp, cho2024ptq4vm, yuan2022ptq4vit, ramachandran2025microscopiq] is a popular method to tackle the high memory and latency demand by mapping full-precision (FP) weights and/or activations to lower-bit representations. In particular, post-training quantization (PTQ) [ramachandran2024clamp, kundu2022bmpq, ramachandran2025accelerating] converts a pre-trained FP model to low-precision and requires calibration on a small dataset [ramachandran2024clamp] to recover model performance. However, PTQ’s calibration data is typically drawn from the original training dataset, potentially restricting its applicability in situations demanding privacy, security of training set [kundu2021analyzing, zhang2023sal].
@@
Consequently, data-free quantization (DFQ), a subset of post-training quantization (PTQ) has emerged as a promising alternative. DFQ allows calibration via generating synthetic data from Gaussian noise [ramachandran2024clamp, KimKK25, li2023psaq, li2022patch] mimicking the distribution of the original train-set. In particular, DFQ for ViTs primarily rely on the distinctive response of self-attention to noise and real data to generate synthetic data [li2023psaq, ramachandran2024clamp]. However, despite their significant progress for ViTs, DFQ techniques for VMM remains largely unexplored. Interestingly, VMM’s recurrent state transition in the S6 layer [gu2023mamba] lacks an explicit self-attention mechanism [ansari2023gpu, ramachandran2023ntrans], limiting the usage of any attention based patch similarity improvement. While the S6 layer can be formulated as a variant of linear attention [han2025demystify], our observations reveal that this implicit attention struggles to distinguish foreground from the background of an image. As observed in LABEL:fig:teaser(c), Fig. 2(a), despite VMM’s “theoretical ability” for global interactions [Azizi2025mambaextend], it fails to capture long-range dependencies in feature representation.
@@
Notably, VMM’s recurrent state transitions introduce dynamic variations in activation characteristics [li2025qmamba]. These variations occur across time steps and tokens [^1] We use token and patch interchangeably based on context., potentially demanding online adaptive outlier selection. This is in sharp contrast to ViTs, which largely demonstrate a static outlier activation pattern [dong2023packqvit]. A contemporary work, QMamba [li2025qmamba], rightly identified the dynamic variations in SSM activations. However, it employs statically determined temporal grouped quantization, unable to adapt to the potential changes in outlier locations, yielding large quantization error at low bit precision. Moreover, QMamba lacks a data-free calibration strategy and limited practical speedup, limiting its use in privacy and latency-sensitive scenarios.
@@
Our Contributions. To address the above challenges in data generation and quantization stages, we present OuroMamba [^2] Inspired by Ouroboros—the self-consuming snake— the name OuroMamba reflects its ability to self-generate data for VMM quantization., a first-of-its-kind method to enable DFQ for VMMs. For data generation, we investigate on VMM’s implicit attention struggles to differentiate foreground with background. In specific, we hypothesize it to be dependent on scanning direction of the compressed hidden state ($h(t)$), which limits explicit spatial token interactions. To mitigate this, we propose OuroMamba-Gen, that essentially enhances the implicit attention through patched spatial interactions in the latent state space, resulting in a refined hidden state namely, patched state $h_{p}(t)$. We then leverage an enhanced implicit attention representation based on $h_{p}(t)$ (LABEL:fig:teaser(d)) to generate synthetic data by employing patch-level contrastive learning [ramachandran2024clamp].
@@
For quantization, we first empirically validate that VMM operation leads to dynamic inter-time-step variations of the outlier channel positions. We then present OuroMamba-Quant, a mixed-precision quantization scheme that performs channel-wise static weight quantization and dynamic activation quantization per-time-step during inference. OuroMamba-Quant reduces the quantization error [lee2025recap] associated to dynamic outliers via three key strategies, (1) a lightweight channel-wise activation outlier detection, (2) mixed-precision quantization, where outlier channels are quantized at higher precision, while remaining inlier channels use lower bit quantization and (3) efficient outlier management using an adaptive outlier list.
[…]
@@ 2 Related Works
Data-Driven PTQ. Existing PTQ techniques for ViTs [lin2021fq, jiang2025adfq, li2023repq, yang2024dopq, ma2024outlier, yuan2022ptq4vit] address long-tailed distributions and static activation outliers but fail to handle VMMs’ dynamically changing outliers [cho2024ptq4vm, li2025qmamba]. While PTQ4VM [cho2024ptq4vm] adapts SmoothQuant [xiao2023smoothquant] to shift activation outliers into weights, it does not quantize SSM activations, limiting its effectiveness. QMamba [li2025qmamba] employs temporal grouped quantization but relies on static groupings, making it unsuitable for ultra-low bit-widths. OuroMamba is the first to identify dynamic outlier variations in VMMs and introduce a mixed-precision quantization framework that achieves SoTA accuracy at low precision.
Date-Free PTQ. Existing DFQ techniques for ViTs, PSAQ-ViT v1 [li2022patch] and v2 [li2023psaq] optimize noise into synthetic data by maximizing global patch similarity entropy, but overlook spatial sensitivity and semantic inter-patch relationships, leading to simplistic data. CLAMP-ViT [ramachandran2024clamp] addresses this by introducing patch-level contrastive learning [ramachandran2024algorithm] on self-attention, aligning foreground patches while separating background patches. Following data generation, these methods apply uniform symmetric quantization to weights and activations, which is suboptimal for VMM. DFQ for VMM thus far remains unexplored, and this work makes the first effort in this direction.
@@ 3.1 Mamba Block in VMM
The S6 Layer. Each Mamba block in VMM utilizes the selective SSM (S6) mechanism [gu2023mamba] that maps an input signal $u(t)\in\mathbb{R}^{E}$ to an intermediate hidden state $h(t)\in\mathbb{R}^{E\times N}$ through a linear recurrent transition. Here, $E$,$N,t$ correspond to the number of channels, state dimension and time-step, respectively. $h(t)$ then produces the output $o(t)\in\mathbb{R}^{E}$ via a linear transformation, as shown below.
@@
where, $S^{+}$ is softplus, $\Delta(t)_{\text{proj}},W_{B},\text{ and }W_{C}$ are linear projection layers. $\Delta(t)\in\mathbb{R}^{E\times 1}$ is the discretization tensor. $\bar{A},\bar{B},C,\Delta$ are input-dependent discrete S6 parameters that exhibit dynamic activation variations at each time step. Eq. 1 deals with timestep-wise inputs. To operate over an input sequence $u$ of batch size $B$, sequence length $M$ with $E$ channels, $u\in\mathbb{R}^{B\times M\times E}$, Eq. 2, Eq. 1 are applied independently to each $u(t)\in\mathbb{R}^{B\times E}$ [gu2023mamba] across $M$ time-steps.
@@
Selective Scan for Vision. While selective scan in S6 is well-suited for 1D NLP tasks, vision data is 2D, requiring VMMs to adopt specialized scanning strategies to flatten images into 1D sequences before processing them through Mamba’s selective scan. Examples include 2-way scanning [zhu2024vision] (Fig. 1(a)), 4-way scanning [liu2025vmamba]. However, scan direction choice can introduce directional dependency, impacting the model’s ability to capture spatial relationships [xiao2024grootvl].
Figure 1: W4A8 quantization performance comparison under the impact of different calibration data sources.
@@ 3.2 Implicit Attention in VMM
Prior works [han2025demystify, ali2024hidden] have shown that S6 layer operation can be reformulated into an implicit self-attention mechanism. By setting $h(0)=0$ and unrolling Eq. 1, the hidden state can be derived as, $h(t)=\sum_{j=1}^{t}(\prod_{k=j+1}^{t}\bar{A}(k))\bar{B}(j)u(j)$ and $o(t)$ can be computed by using this derived value of $h(t)$ in Eq. 1. Additionally, in computing $o(t)$, the contribution of the $j^{th}$ token on the $i^{th}$ token is captured as follows,
$\small\tilde{\alpha}[i,j]=\sum_{m=1}^{N}\alpha^{m}[i,j]=\sum_{m=1}^{N}C(i)(\prod_{k=j+1}^{i}\bar{A}(k)_{m})\bar{B}(j)[m]$ (4)
@@ 3.3 Quantization
In this paper, we perform symmetric group quantization [ramachandran2024clamp, li2022patch] of both weights and activations for VMMs. The quantization process for an input FP tensor $X$ with a target bit-precision $b$ is given by,
@@ 3.4 Contrastive Objective
Contrastive learning based on the infoNCE loss [wang2023positive] helps learn an anchor patch from both similar (positive) and dissimilar (negative) patches. Following [ramachandran2024clamp], the formulation of patch-level contrastive loss ($\mathcal{L}^{C}_{l,(i,j)}$) we employ for data generation is as follows,
@@
where, *, ++, and - correspond to the anchor patch, positive, and negative patches, respectively. $\tau$ controls the concentration level [wang2023positive]. $\lambda_{l,(i,j)}$ (shown without subscript above for brevity) represents the embedding employed to calculate $\mathcal{L}^{C}_{l,(i,j)}$. $l$ is the layer and $(i,j)$ is the location of anchor patch. The final loss is given as $\mathcal{L}^{C}=\sum_{l}\sum_{i}\sum_{j}\mathcal{L}^{C}_{l,(i,j)}$.
@@ 4 Motivational Analysis
Figure 2: (a) Forward and backward SSM state transition of first layer of Vim-S [zhu2024vision], (b) Naive synthetic data samples generated by applying [ramachandran2024clamp] on VMM implicit attention.
@@ 4.1 Synthetic Data Generation
We investigate the W4A8 quantization performance of Vim-S [zhu2024vision] on an existing VMM PTQ framework, QMamba [li2025qmamba], using different calibration data sources. In specific, we perform calibration with both real data and ViT generated synthetic data of same sample size of 1024 images. We use synthetic calibration data produced by ViT DFQ technique, namely, CLAMP-ViT [ramachandran2024clamp]. Fig. 1 reveals that across tasks, ViT synthetic data results in significant accuracy degradation (7-14%) as compared to using real calibration data.
@@
As shown in LABEL:fig:teaser(c), VMM’s implicit attention struggles to distinguish foreground from background. This is potentially due to the forgetfulness of the states [Azizi2025mambaextend] as we traverse along the scanning direction (Fig. 2(a)) [han2025demystify]. This, unlike self-attention, limits long-range global interactions. Additionally, VMMs [zhu2024vision, liu2025vmamba] employ multiple scanning directions (Sec. 3.1) to improve global interactions. However, they process each direction independently with separate S6 blocks. Consequently, each scanning direction’s implicit attention exhibits directional bias and are incoherent [han2025demystify].
Following [ramachandran2024clamp], that applies patch-level contrastive loss to ViT self-attention to generate synthetic data, we apply the same method to VMM’s forward scan S6 block and present the generated data in Fig. 2(b). The “naive synthetic data” appears noisy, lacks semantic structure, and performs worse than ViT-generated synthetic data for calibration, Fig. 1). Extending this to other scanning directions yields similar results (see Sec. 6). These observations motivate an optimal VMM data generation technique, capturing spatial dependencies without scanning direction constraints.